# Efficient and Reliable Hitting-Set Computations for the Implicit Hitting Set Approach

### v1

@@ 7 $\gamma\leftarrow\textnormal{{Optimize}}(\mathcal{K},O,ub,\text{opt?})$;
Algorithm 2 details a generalized abstraction of the Solve-HS procedure (the hitting set component), allowing for different ways of combining IP solving, conflict-driven pseudo-Boolean solving, and stochastic local search. Assume PBO-IHS (Algorithm 1) is tasked with computing an optimal solution to an instance $(F,O)$. It will invoke Solve-HS with the set $\mathcal{K}$ of cores extracted so far, the objective $O$, and the current upper and lower bound $(ub,lb)$ on the optimal cost of $(F,O)$.
First Solve-HS invokes use-sls, a user-specified heuristic for determining whether SLS should be invoked on Line 2. If the solution $\gamma$ that SLS returns has a cost lower than the current upper bound $ub$, Solve-HS terminates on Line 2. Otherwise, the subroutine Optimize is invoked on Line 2 and tasked to either compute a solution $\gamma$ of $(\mathcal{K},O)$ that has $\gamma(O)<ub$, or (if no such solutions exist) an optimal solution $\gamma$ of $(\mathcal{K},O)$. Since the current best solution $\alpha_{\mathit{best}}$ of $(F,O)$ maintained by PBO-IHS has $\alpha_{\mathit{best}}(O)=ub$ and is also a solution of $(\mathcal{K},O)$, an alternative view on Optimize is that of finding a solution of $(\mathcal{K},O)$ of cost lower than $ub$, or determining that one does not exist. As an additional heuristic, we allow forcing Optimize to compute an optimal solution to $(\mathcal{K},O)$, essentially forcing an improvement to the lower bound. More specifically, the user-specified heuristic optimal-sol invoked on Line 2 returns true if Optimize is only allowed to terminate with an optimal solution to $(\mathcal{K},O)$.
@@
Next, we detail the different parts of Solve-HS the different instantiations of Optimize that we consider as well as integration of local search for potential practical runtime improvements. Our general view on computing hitting sets allows for combining these instantiations in various different ways.
@@
The arguably most commonly used instantiation of Optimize invokes a commercial MIP solver on $(\mathcal{K},O)$. We consider two variants of IP solving for instantiating Optimize. The first is IP solving techniques that make use of fast floating-point arithmetic. These are arguably among the most efficient methods of computing solutions $(\mathcal{K},O)$, and most state-of-the-art IHS approaches instantiate Solve-HS with a floating-point IP solver. One disadvantage of such solvers, however, is that there have been repeated reports of such solvers outputting faulty solutions [CKSW13], preventing their use in applications that require trust in the results. Thus, we also consider exact IP solving, i.e., IP solving methods that employ exact arithmetic to circumvent issues related to floating point precision and rounding errors [CKSW13, EG23]. This increased reliability of the results does, however, come at a cost in terms of computational efficiency. Recent computational studies [EG23] report slowdowns of a factor $8.1$ when comparing a state-of-the-art exact MIP solver to its floating point counterpart.
@@
Complementing the Branch & Cut search commonly employed by IP solvers, we also consider harnessing recent developments in native pseudo-Boolean optimization procedures that reduce the problem of computing an optimal solution of $(\mathcal{K},O)$ into a sequence of decision problems, which are tackled with a decision procedure similar to the one employed within Algorithm 1 to extract cores [DGD+21, DGN21]. More specifically, we consider three different variants of pseudo-Boolean optimization algorithms for computing an optimal solution to $(\mathcal{K},O)$: solution-improving search (SIS), core-guided search (CG), and core-boosted search (CB).
@@
A SIS-based approach maintains an upper bound $ub_{c}$ on the optimal cost of $(\mathcal{K},O)$, In the context of Solve-HS, $ub_{c}$ is instantiated to the input upper bound $ub$. On each iteration, the decision procedure is invoked on $\mathcal{K}\land(O<ub_{c})$, i.e., the cores in $\mathcal{K}$ and a solution improving constraint $(O<ub_{c})$. If a solution exists, the returned solution $\gamma$ will have $\gamma(O)<ub_{c}$, so the upper bound is updated, and the process repeated. Otherwise (if the decision procedure reports unsatisfiable), the latest solution found is determined to be optimal for $(\mathcal{K},O)$. In addition to the basic SIS approach just described, we consider a variant that we call reified SIS. An incremental SIS approach adds the reified solution-improving constraint $r\Rightarrow(\gamma(O)<ub_{c})$, i.e., the linear inequality $M\overline{r}+\gamma(O)<ub_{c}$ where $r$ is a fresh variable and $M$ is a large enough constant, and invokes the decision procedure while assuming $r=1$. The benefit of relaxing the solution improving constraint is that the decision procedure instantiation can be incrementally maintained between iterations of Solve-HS.
@@
A core-guided approach [DGD+21] maintains a reformulated objective $O^{R}$, initialized to $O$, and a set $\mathcal{C}$ of additional constraints, initialized to $\emptyset$. On each iteration, the decision procedure is queried for a solution of $\mathcal{K}\land\mathcal{C}$ while assuming all literals in $O^{R}$ to 0. More informally, the query asks for a solution $\gamma$ of $\mathcal{K}\land\mathcal{C}$ that has $\gamma(O^{R})=0$. If a solution exists, the obtained solution is an optimal solution of $(\mathcal{K},O)$ and the procedure terminates. Otherwise, the call returns a new core falsified by the assumptions. Next, the core-guided approach introduces new constraints to $\mathcal{C}$ and reformulates $O^{R}$ in a manner that resolves the inconsistency represented by the core in a cost-minimal way. Informally speaking, the invariant maintained by a core-guided search is that optimal solutions of $(\mathcal{K}\land\mathcal{C},O^{R})$ are optimal solutions of $(\mathcal{K},O)$. Since the set $\mathcal{K}$ of cores that we invoke core-guided search on monotonically increases during search, the state of the core-guided solver is maintained between invocations. While such incremental invocations of core-guided search prevent fixing any variables based on the bounds on the optimal cost of $(\mathcal{K},O)$, variables are still fixed based on the bounds of the full instance $(F,O)$. For a concrete example, we fix the value of an objective literal $\ell$ that has a constant larger than $ub$ to 0 since any solution that assigns $\ell=1$ would incur a cost larger than ubub
@@
Finally, core-boosted search [BDS19] is a hybrid of core-guided and solution-improving search that starts with core-guided search for a user-specified number of iterations. Then it invokes solution-improving search on the final reformulated instance $(\mathcal{K}\land\mathcal{C},O^{R})$ obtained by core-guided search.
@@
It is important to note that SIS, CG and CB can each provide intermediate solutions that allow terminating search as soon as a solution of cost lower than $ub$ is found. Specifically, the instantiation of core-guided search that we use makes use of the so-called stratification heuristic [ABGL13] that results in it obtaining intermediate solutions, even if pure core-guided search only ever finds an optimal solution.
@@ 0.3.2 Integrating Stochastic Local Search
Stochastic local search (SLS) is a lightweight heuristic approach to pseudo-Boolean optimization that quickly finds low-cost solutions. In contrast to the approaches described in Section 0.3.1, SLS alone can not give any guarantee on optimality of a solution and cannot therefore establish a lower bound on its own. Hence an IHS approach with an SLS integration in the hitting set component cannot terminate without an instantiation of Optimize that can check whether the solution found by SLS is optimal.
@@
In more detail, we design Core-SLS for incremental calls, to run in conjunction with another instantiation of Optimize, and to produce good solutions for core extraction. Adapting the SLS approach to incremental calls is straightforward, as we can keep the state from the previous call and simply add the new constraints representing additional cores. Effectively, Core-SLS then tries to repair the previous solution to satisfy the new constraints. and (iii) how flipping the variable value would impact the cost of the assignment.
The solution returned by Solve-HS is used to extract the next set of cores. Thus it is important that the solutions returned by Solve-HS are not only of low cost, but guide the IHS solver towards an optimal solution. We incorporate this idea in our SLS integration by ensuring that the solutions are diversified, i.e., that the subsequent solutions returned by Solve-HS have a large Hamming-distance, whenever possible. Hence, we split Core-SLS into two phases: (i) we start from the previous solution and (ii) we restart the search and start from the previous solution but randomly flip a large number of variable values. Whenever we find a solution in both phases, we prefer the one of lower cost and in case of a tie, we prefer the solution from the second phase to foster diversification. Further, Core-SLS’s purpose is avoiding costly calls to the optimizer. Hence, we skip Core-SLS whenever we know from prior iterations that the optimizer calls do not take longer than Core-SLS calls.
@@ 0.4 Certified IHS Computations
We now turn our attention to developing proofs for the IHS paradigm, i.e., independently checkable certificates of correctness of the result. The VeriPB format is a general-purpose proof format for pseudo-Boolean optimization, combining implicational reasoning using cutting planes operations (such as adding up two PB constraints) [CCT87] with well-chosen strengthening rules that allow deriving non-implied constraints [BGMN23]. For representational succinctness of the current paper, details of the actual proof rules are not critical to spell out, but it is relevant that the strengthening rules allow us to derive fresh reification variables. In other words, if $x$ is a variable that does not yet appear in the proof and $C$ is a PB constraint, we can always derive (one or both) two new PB constraints $x\Rightarrow C$ and $x\Leftarrow C$ that express that $x$ is true precisely when $C$ is satisfied. If $C$ is the normalized constraint ${\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$, then $x\Rightarrow C$ and $x\Leftarrow C$ are $A\overline{x}+{\textstyle\sum}_{i}a_{i}\ell_{i}\geq A\text{ and }({\textstyle\sum}_{i}a_{i}-A+1)x+{\textstyle\sum}_{i}a_{i}\overline{\ell}_{i}\geq{\textstyle\sum}_{i}a_{i}-A+1\text{, respectively. }$
@@ 0.4.1 VeriPB Proofs for PBO-IHS
As instantiations of the conflict-driven pseudo-Boolean approach discussed, e.g., in Section 0.3.1 have had support for VeriPB proof logging for a while (see, e.g., the recent certifying track in the PBO Competitions [Rou24]), a natural idea toward obtaining uniform proofs for PBO-IHS is to instantiate Solve-HS with a PB-based approach and to combine the proofs produced by the Solve-HS subroutine with the proofs produced by the Extract-Cores subroutine. While the idea is conceptually simple, there are a number of details that need to be considered. Due to space restrictions, we cannot go into detail about how the proof production for the different components works. Instead, we discuss challenges in their integration and how they are addressed.
@@
With two different solvers co-producing parts of the proof, neither of the solvers are fully aware of which constraints are derived in the proofs and (in case they introduce new variables), which variables are unused so far. To facilitate this, we developed a general pseudo-Boolean proof logging API which both solvers use and which is responsible for maintaining relevant information, such as mappings from solver-specific variables to their name used in the proof. We expect this API to be useful more generally for future pseudo-Boolean proof logging approaches as well.
@@
Another issue to be dealt with is that the optimizer uses (Solve-HS) is not optimizing the original problem, but a derived problem. When using a pseudo-Boolean optimizer with proof logging, it will log solutions it finds and might then use the associated solution-improving constraint (which expresses that from now on we are only interested in better solutions). Taking a holistic perspective, however, this solution-improving constraint is not sound: the solution Solve-HS finds is not a solution of the whole problem. Above we already described two versions of integrating a SIS solver (incremental and non-incremental). In the incremental case the solver does not introduce a proper solution-improving constraint but a reified version thereof; this is easy to deal with from a proof logging perspective. In the non-incremental version, however, the solver is ran without assumptions and restarted upon every call. In this case, the solver will be deriving those unsound solution-improving constraints. Our solution for dealing with this is to allow for a small discrepancy between the constraints the solver sees and the constraints derived in the proof, making sure the solution-improving constraint logged in the proof is indeed reified.

### v2

@@ 7 $\gamma\leftarrow\textnormal{{Optimize}}(\mathcal{K},O,ub,\text{opt?})$;
Algorithm 2 details a generalized abstraction of the Solve-HS procedure (the hitting set component), allowing for different ways of combining IP solving, conflict-driven pseudo-Boolean solving, and stochastic local search. PBO-IHS (Algorithm 1) is tasked with computing an optimal solution to an instance $(F,O)$. It will invoke Solve-HS with the set $\mathcal{K}$ of cores extracted so far, the objective $O$, and the current upper and lower bound $(ub,lb)$ on the optimal cost of $(F,O)$.
First Solve-HS invokes use-sls, a user-specified heuristic for determining whether SLS should be invoked on Line 2. If the solution $\gamma$ that SLS returns has a cost lower than the current upper bound $ub$, Solve-HS terminates on Line 2. Otherwise, the subroutine Optimize is invoked on Line 2 and tasked to either compute a solution $\gamma$ of $(\mathcal{K},O)$ that has $\gamma(O)<ub$, or (if no such solutions exist) an optimal solution $\gamma$ of $(\mathcal{K},O)$. Since the current best solution $\alpha_{\mathit{best}}$ of $(F,O)$ maintained by PBO-IHS has $\alpha_{\mathit{best}}(O)=ub$ and is also a solution of $(\mathcal{K},O)$, an alternative view on Optimize is that of either finding a solution of $(\mathcal{K},O)$ of cost lower than $ub$ or determining that such a solution does not exist. As an additional heuristic, we allow forcing Optimize to compute an optimal solution to $(\mathcal{K},O)$, essentially forcing an improvement to the lower bound. More specifically, the user-specified heuristic optimal-sol invoked on Line 2 returns true if Optimize is only allowed to terminate with an optimal solution to $(\mathcal{K},O)$.
@@
Next, we detail different instantiations of Optimize that we consider as well as integration of local search for potential practical runtime improvements. Our general view on computing hitting sets allows for combining these instantiations in various different ways.
@@
The arguably most commonly used instantiation of Optimize invokes a commercial IP solver on $(\mathcal{K},O)$. We consider two variants of IP solving for instantiating Optimize. The first variant employs fast floating-point arithmetic in its computation. These are arguably among the most efficient methods of computing solutions $(\mathcal{K},O)$, and most state-of-the-art IHS approaches instantiate Solve-HS with a floating-point IP solver. One disadvantage of such solvers, however, is that there have been repeated reports of such solvers outputting faulty solutions [CKSW13], preventing their use in applications that require trust in the results. Thus, we also consider exact IP solving, i.e., IP solving methods that employ exact arithmetic to circumvent issues related to floating point precision and rounding errors [CKSW13, EG23]. This increased reliability of the results does, however, come at a cost in terms of computational efficiency. Recent computational studies [EG23] report slowdowns of a factor $8.1$ when comparing a state-of-the-art exact IP solver to its floating point counterpart.
@@
Complementing the branch & cut search commonly employed in IP solvers, we also consider harnessing recent developments in native pseudo-Boolean optimization procedures that reduce the problem of computing an optimal solution of $(\mathcal{K},O)$ into a sequence of decision problems. The decision problems are tackled with a decision procedure similar to the one employed within Algorithm 1 to extract cores [DGD+21, DGN21]. More specifically, we consider three different variants of pseudo-Boolean optimization algorithms for computing an optimal solution to $(\mathcal{K},O)$: solution-improving search (SIS), core-guided search (CG), and core-boosted search (CB).
@@
A SIS-based approach maintains an upper bound $ub_{c}$ on the optimal cost of $(\mathcal{K},O)$, In the context of Solve-HS, $ub_{c}$ is instantiated to the input upper bound $ub$. In each iteration, the decision procedure is invoked on $\mathcal{K}\land(O<ub_{c})$, i.e., the cores in $\mathcal{K}$ and a solution-improving constraint $(O<ub_{c})$. If a solution exists, the returned solution $\gamma$ will have $\gamma(O)<ub_{c}$, so the upper bound is updated, and the process repeats. Otherwise (if the decision procedure reports unsatisfiable), the latest solution found is determined to be optimal for $(\mathcal{K},O)$.
One potential issue with this SIS-based approach is that the constraints $O<ub_{c}$ that are added to the solver are not sound for future calls to this oracle. Indeed, when new cores are learned, $ub_{c}$ might no longer be an upper bound on the best objective value. As a consequence, a fresh instantiation of the optimization solver needs to be created each time it is called. To deal with this, we also consider a variant that we call reified SIS. An incremental SIS approach adds the reified solution-improving constraint
$r_{ub_{c}}\Rightarrow(\gamma(O)<ub_{c}),$
i.e., the linear inequality
$M\overline{r_{ub_{c}}}+\gamma(O)<ub_{c}$
where $r_{ub_{c}}$ is a fresh variable and $M$ is a large enough constant, and invokes the decision procedure while assuming $r=1$. The benefit of relaxing the solution-improving constraint is that the decision procedure instantiation can be incrementally maintained between iterations of Solve-HS. Not just the constraint itself, but also all constraints derived from it will be reactivated in all future solve calls where a bound of $ub_{c}$ or lower is enforced.
@@
A core-guided approach [DGD+21] maintains a reformulated objective $O^{R}$, initialized to $O$, and a set $\mathcal{C}$ of additional constraints, initialized to $\emptyset$. In each iteration, the decision procedure is queried for a solution of $\mathcal{K}\land\mathcal{C}$ while assuming all literals in $O^{R}$ to 0. More informally, the query asks for a solution $\gamma$ of $\mathcal{K}\land\mathcal{C}$ that has $\gamma(O^{R})=0$. If a solution exists, the obtained solution is an optimal solution of $(\mathcal{K},O)$ and the procedure terminates. Otherwise, the call returns a new core falsified by the assumptions. Next, the core-guided approach introduces new constraints to $\mathcal{C}$ and reformulates $O^{R}$ in a manner that resolves the inconsistency represented by the core in a cost-minimal way. Informally speaking, the invariant maintained by a core-guided search is that optimal solutions of $(\mathcal{K}\land\mathcal{C},O^{R})$ are optimal solutions of $(\mathcal{K},O)$. Since the set $\mathcal{K}$ of cores that we invoke core-guided search on monotonically increases during search, the state of the core-guided solver is maintained between invocations. While such incremental invocations of core-guided search prevent fixing any variables based on the bounds on the optimal cost of $(\mathcal{K},O)$, variables are still fixed based on the bounds of the full instance $(F,O)$. For a concrete example, we fix the value of an objective literal $\ell$ that has a constant larger than $ub$ to 0 since any solution that assigns $\ell=1$ would incur a cost larger than ubub
@@
Finally, core-boosted search [BDS19] is a hybrid of core-guided and solution-improving search that starts with core-guided search for a user-specified number of iterations. After this, solution-improving search is invoked on the final reformulated instance $(\mathcal{K}\land\mathcal{C},O^{R})$ obtained from the core-guided search phase.
@@
It is important to note that SIS, CG and CB can each provide intermediate solutions that allow terminating search as soon as a solution of cost lower than $ub$ is found. Specifically, the instantiation of core-guided search that we use makes use of the so-called stratification heuristic [ABGL13] and weight-aware core extraction [BJ17] that provide intermediate solutions, whereas plain core-guided search does not find any intermediate solutions before finding the first solution which is guaranteed to be optimal.
@@ 0.3.2 Integrating Stochastic Local Search
Stochastic local search (SLS) is a lightweight heuristic approach to pseudo-Boolean optimization that quickly finds low-cost solutions. In contrast to the approaches described in Section 0.3.1, SLS alone can not give any guarantee on optimality of a solution and cannot therefore establish a lower bound on its own. Hence an IHS approach with SLS integration in the hitting set component cannot terminate without an instantiation of Optimize that can check whether the solution found by SLS is optimal.
@@
In more detail, we design Core-SLS for incremental calls, to run in conjunction with another instantiation of Optimize, and to produce good solutions for core extraction. Adapting the SLS approach to incremental calls is straightforward, as we can keep the state from the previous call and simply add the new constraints representing additional cores. Effectively, Core-SLS seeks to repair the previous solution to satisfy the new constraints. and (iii) how flipping the variable value would impact the cost of the assignment.
The solution returned by Solve-HS is used to extract the next set of cores. Thus it is important that the solutions returned by Solve-HS are not only of low cost, but also guide the IHS solver towards an optimal solution. We incorporate this idea in our SLS integration by ensuring that the solutions are diversified, i.e., that the subsequent solutions returned by Solve-HS have a large Hamming-distance, whenever possible. Hence, we split Core-SLS into two phases: (i) we start from the previous solution and (ii) we restart the search and start from the previous solution but randomly flip a large number of variable values. Whenever we find a solution in both phases, we prefer the one of lower cost. In case of a tie, we prefer the solution from the second phase to foster diversification. Further, Core-SLS’s purpose is avoiding costly calls to the optimizer. Hence we skip Core-SLS whenever we know from prior iterations that the optimizer calls do not take longer than Core-SLS calls.
@@ 0.4 Certified IHS Computations
We now turn to developing proofs, i.e., independently checkable certificates of correctness of results, for the IHS paradigm. We employ the VeriPB format as a general-purpose proof format for pseudo-Boolean optimization, combining implicational reasoning using cutting planes operations (such as adding up two PB constraints) [CCT87] with well-chosen strengthening rules that allow deriving non-implied constraints [BGMN23]. For representational succinctness of the current paper, details of the actual proof rules are not critical to spell out, but it is relevant that the strengthening rules allow us to derive fresh reification variables. In other words, if $x$ is a variable that does not yet appear in the proof and $C$ is a PB constraint, we can always derive (one or both) two new PB constraints $x\Rightarrow C$ and $x\Leftarrow C$ that express that $x$ is true precisely when $C$ is satisfied. If $C$ is the normalized constraint ${\textstyle\sum}_{i}a_{i}\ell_{i}\geq A$, then $x\Rightarrow C$ and $x\Leftarrow C$ are
$A\overline{x}+{\textstyle\sum}_{i}a_{i}\ell_{i}\geq A\text{ and }$
$({\textstyle\sum}_{i}a_{i}-A+1)x+{\textstyle\sum}_{i}a_{i}\overline{\ell}_{i}\geq{\textstyle\sum}_{i}a_{i}-A+1,$
respectively.
@@ 0.4.1 VeriPB Proofs for PBO-IHS
Pseudo-Boolean solvers using solution-improving search and core guided search have had support for VeriPB proof logging for a while (see, e.g., the recent certifying track in the PBO Competitions [Rou24]). As such, a natural idea toward obtaining uniform proofs for PBO-IHS is to instantiate Solve-HS with a PB-based approach and to combine the proofs produced by the Solve-HS subroutine with the proofs produced by the Extract-Cores subroutine. While the idea is conceptually simple, there are a number of details that need to be considered. Due to space restrictions, we do not go into detail about how the proof production for the different components works. Instead, we discuss challenges in their integration and how they are addressed.
@@
With two different solvers co-producing parts of the proof, neither of the solvers are fully aware of which constraints are derived in the proofs and, in case they introduce new variables, which variables are unused so far. To facilitate this, we developed a general pseudo-Boolean proof logging API which both solvers use and which is responsible for maintaining relevant information, such as mappings from solver-specific variables to their name used in the proof. We expect this API to be useful more generally for future pseudo-Boolean proof logging approaches as well.
@@
Another issue that needs to be addressed is that the optimizer uses (Solve-HS) is not optimizing the original problem, but a derived problem. When using a pseudo-Boolean optimizer with proof logging, it will log solutions it finds and might then use the associated solution-improving constraint expresses that from now on we are only interested in better solutions. Taking a holistic perspective, however, this solution-improving constraint is not sound: the solution Solve-HS finds is not a solution of the whole problem. We already described in Section 3 two versions of integrating a SIS solver (incremental and non-incremental). In the incremental case the solver does not introduce a proper solution-improving constraint but a reified version thereof; this is easy to deal with from a proof logging perspective. In the non-incremental version, however, the solver is ran without assumptions and restarted upon every call. In this case the solver will be deriving those unsound solution-improving constraints. Our solution for dealing with this is to allow for a small discrepancy between the constraints the solver sees and the constraints derived in the proof, ensuring that the solution-improving constraint logged in the proof is indeed reified.
Example 2.
Assume that the objective at hand is $x_{1}+2x_{2}+3x_{3}$ and during an execution of Solve-HS we find a solution with $x_{1}=x_{2}=1$ and $x_{3}=0$. While this is a solution of $\mathcal{K}$, we do not know if this can be extended to a solution of $F$. Deriving the solution-improving constraint
$x_{1}+2x_{2}+3x_{3}$ $<3$ (1)
$\text{(which normalizes to }\overline{x}_{1}+2\overline{x}_{2}+3\overline{x}_{3}$ $\geq 4\text{)}$ (2)
in the proof would be unsound. In reified SIS this is not a problem since this constraint is not learned unconditionally. In regular SIS, however, when adding proof logging, we do not wish to change the behaviour of the solver. To this end, we add the above constraint to the solver but in the proof instead derive the reification constraint
$4\overline{r}_{3}+\overline{x}_{1}+2\overline{x}_{2}+3\overline{x}_{3}$ $\geq 4$ (3)
Later on, whenever the solver uses constraint (2) for deriving new constraints internally, the proof will instead reference constraint (3), resulting in all derivations of it to be conditional on $r_{3}$. If it later turns out that there is solution of cost $3$ for the original PBO instance, we can derive without loss of generality that $r_{3}$ is indeed true.